OpenStudy (anonymous):

Find the polynomial function with roots 1, –2, and 5. help me please

OpenStudy (anonymous):

start with \[(x-1)(x+2)(x-5)\] and multiply out

OpenStudy (anonymous):

(x-1)(x+2)(x-5) =(x^2+x-2)(x-5) =x^3+x^2-2x-5x^2-5x+10 =x^3-4x^2-7x+10

OpenStudy (ash2326):

if we are given roots say a, b and c. The polynomial will be (x-a)(x-b)(x-c) here the roots are 1, -2 and 5 (x-1)(x+2)(x-5) \[(x^2+x-2)(x-5)\] \[x^3+x^2-2x-5x^2-5x+10\] combine like terms \[x^3-4x^2-7x+10\]

OpenStudy (anonymous):

Thank you so much you guys!<3